# Learning to Instruct for Visual Instruction Tuning

### v1

@@ 4 Related Work
Visual Instruction Tuning. The concept of Visual Instruction Tuning was first introduced in LLaVA (Liu et al., 2023a) and MiniGPT-4 (Zhu et al., 2024), aiming to unify the understanding of vision and language by leveraging pre-trained visual and language models. Common system architectures typically consist of (1) a pre-trained visual model for encoding visual features, (2) a pre-trained large language model for interpreting images and user instructions and generating responses, and (3) a cross-modal connector for aligning visual features with the language model’s input. Visual resamplers, such as Qformer (Li et al., 2023a), can serve as an optional module to reduce the number of visual patches (Bai et al., 2023a; Dai et al., 2023). LLaVA-NeXT (Liu et al., 2024b) significantly enhances visual perception by using dynamic visual resolutions. DEEM (Luo et al., 2024) replaces the traditional visual encoder with a diffusion model, further enhancing visual perception. Cambrain-1 (Tong et al., 2024) improves visual robustness through visual encoder routing, but it also introduces higher training overhead.
@@
The most relevant recent work to ours is IM (Shi et al., 2024), which incorporates loss over instructions during the instruction tuning process of language models. However, our work differs significantly in the following key aspects: 1. Scope: Our work focuses on MLLMs, whereas IM is on language models. 2. Motivation: Our primary motivation is to encourage the model to pay greater attention to visual information and avoid shortcut learning based solely on textual cues. In contrast, IM is primarily aimed at mitigating overfitting. 3. Methodological Differences: We employ an automated approach to filter out the impact of frequently occurring template instructions (see ablation study in Table 3), whereas IM only filters a limited set of special tokens, such as “<|user|>”. 4. Scalability: We validate the effectiveness of LIT on a large-scale dataset of nearly 1M samples, whereas IM’s evaluation is limited to much smaller datasets, showing performance saturation at just 13k samples.
@@ 5 Conclusion
In this paper, we propose LIT, which enhances multimodal capabilities by learning to instruct images as a regularizer, rather than focusing only on learning to respond. This enables LIT to seamlessly expand the training data, reducing overfitting and proactively guiding the model to learn visual inputs more effectively, thereby avoiding shortcuts. Experimental results demonstrate the effectiveness of LIT across 16 multimodal tasks, highlighting its superior performance on OCR and image captioning tasks by placing greater emphasis on visual content. Furthermore, LIT significantly improves hallucination mitigation. It is also worth noting that our method is orthogonal to existing advancements in MLLMs and can be easily integrated into these methods with minimal compromises on computational costs. We believe that LIT has the potential to evolve the general VIT framework by mitigating overfitting, enhancing data efficiency, and promoting improved generalization in MLLMs.
Impact Statement
@@ Appendix A Detailed Removed Task Templates
In this section, we provide a comprehensive overview of the task templates removed during the training of TinyLLaVA, LLaVA 1.5, and LLaVA-NeXT, as shown in Table 5 and Table 6. Notably, LLaVA-NeXT-Data includes a more diverse set of instruction datasets, leading to a wider variety of task templates. Consequently, a larger number of task templates are removed compared to those discarded from the LLaVA-mix-665k dataset.
@@ B.1 More Details on Hyperparameter
Detailed hyperparameters for training TinyLLaVA, LLaVA 1.5 and LLaVA-NeXT are shown in Table 7. For TinyLLaVA and LLaVA 1.5, pretraining is conducted with a learning rate of 1e-3 and a batch size of 256, while finetuning uses a learning rate of 2e-5 and a batch size of 128. For LLaVA-NeXT, pretraining uses a learning rate of 1e-3 and a batch size of 128, with finetuning using a learning rate of 1e-5 and a batch size of 32. All experiments use the AdamW optimizer (Loshchilov & Hutter, 2019) and a cosine decay schedule with a warm up ratio of 0.03. During pretraining phase, only the multimodal connector is trained. In finetuning, TinyLLaVA and LLaVA 1.5 jointly train the connector and language model, while LLaVA-NeXT trains all parameters, including the vision encoder.
@@ B.2 More Details on Training Data
In the finetuning stage of instruction tuning, we use the LLaVA-mix-665k data (Liu et al., 2024a) on TinyLLaVA and LLaVA 1.5, which incorporates a diverse set of instruction-following datasets, including free conversational data (LLaVA-Instruct (Liu et al., 2023a)), visual question answering (VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), OKVQA (Marino et al., 2019), A-OKVQA (Schwenk et al., 2022)), OCR and captioning (OCRVQA (Mishra et al., 2019), TextCaps (Sidorov et al., 2020)), and visual grounding (RefCOCO (Kazemzadeh et al., 2014), VG (Krishna et al., 2017)). For LLaVA-NeXT, we use the LLaVA-NeXT-Data (Liu et al., 2024b), which extends the LLaVA-mix-665k (Liu et al., 2024a) dataset with high-quality user instruction data from LAION-GPT-V, ShareGPT-4V (Chen et al., 2024a), and LLaVA-demo (Liu et al., 2023a), as well as OCR, document, and chart data from DocVQA (Mathew et al., 2021), SynDog-EN (Kim et al., 2022), ChartQA (Masry et al., 2022), DVQA (Kafle et al., 2018), and AI2D (Kembhavi et al., 2016).
@@ Appendix C More Details on CHAIR Evaluation
As mention in Section 3.3, we utilize the CHAIR metric to evaluate object hallucination in image captioning tasks. The metric assesses object hallucinations through two dimensions: per-instance $\text{CHAIR}_{i}$ and per-sentence $\text{CHAIR}_{s}$. The former measures the fraction of object instances that are hallucinated, while the latter determines the proportion of sentences that contain at least one hallucinated object. Definitions of these two metrics are formally provided:
@@
CHAIR =${}_{s}=\frac{|\{\text{sentences with hallucinated object}\}|}{|\{\text{ all sentences}\}|}.$
We follow (Huang et al., 2023) to perform the CHAIR evaluation on the COCO2014 dataset. Specifically, we randomly select 500 images in the validation set. The decoding process uses the greedy strategy and the Beam search with $N_{beam}=5$, with a maximum of 512 tokens for new content generated.
@@ Appendix D More Cases
More cases of our LIT compared to VIT are shown in Figure 10. Our LIT consistently outperforms VIT across a range of scenarios, including OCR, image captioning, and hallucination mitigation.
Figure 10: More cases of LIT and VIT across OCR, image captioning, and hallucination mitigation tasks.
@@ Appendix E Computational Cost
Table 8 presents the computational cost of VIT and LIT across different instruction-to-response length ratios. The experiments are conducted on 8 NVIDIA A100 GPUs.
Table 8: Computational cost of VIT and LIT across different instruction-to-response length ratios. We report the number of samples per batch and steps per batch. Experiments are based on LLaVA-1.5 Vicuna-7B.
@@
LIT 2.640 0.330 2.623 0.328 2.676 0.334 2.613 0.327 2.708 0.338
Appendix F Limitations
Despite the promising capabilities underscored by LIT, several limitations must be acknowledged. First, the performance improvements achieved by LIT mainly stem from the informative instruction content that aids in better understanding visual inputs. In contrast, redundant instruction content, such as system or task templates, do not contribute to performance gains, as they lack relevant information about the visual content. Second, it is critical to ensure that the instructions do not contain harmful or biased content. The presence of such content could raise concerns about the fairness and reliability of MLLMs, making them more susceptible to generating flawed, biased, or even harmful responses.

### v2

@@ 4 Related Work
Visual Instruction Tuning. Recent advances in multimodal learning (Radford et al., 2021; Zhou et al., 2025; Zhao et al., 2025; Peng et al., 2025a, b; Liu et al., 2025) have greatly enhanced the integration of visual and linguistic understanding. Building upon these foundations, the concept of Visual Instruction Tuning was first introduced in LLaVA (Liu et al., 2023a) and MiniGPT-4 (Zhu et al., 2024), aiming to unify the understanding of vision and language by leveraging pre-trained visual and language models. Common system architectures typically consist of (1) a pre-trained visual model for encoding visual features, (2) a pre-trained large language model for interpreting images and user instructions and generating responses, and (3) a cross-modal connector for aligning visual features with the language model’s input. Visual resamplers, such as Qformer (Li et al., 2023a), can serve as an optional module to reduce the number of visual patches (Bai et al., 2023a; Dai et al., 2023). LLaVA-NeXT (Liu et al., 2024b) significantly enhances visual perception by using dynamic visual resolutions. DEEM (Luo et al., 2024) replaces the traditional visual encoder with a diffusion model, further enhancing visual perception. Cambrain-1 (Tong et al., 2024) improves visual robustness through visual encoder routing, but it also introduces higher training overhead.
@@
The most relevant recent work to ours is IM (Shi et al., 2024), which incorporates loss over instructions during the instruction tuning process of language models. However, our work differs significantly in the following key aspects: 1. Scope: Our work focuses on MLLMs, whereas IM is on language models. 2. Motivation: Our primary motivation is to encourage the model to pay greater attention to visual information and avoid shortcut learning based solely on textual cues. In contrast, IM is primarily aimed at mitigating overfitting. 3. Methodological Differences: We employ an automated approach to filter out the impact of frequently occurring template instructions (see ablation study in Table˜4), whereas IM only filters a limited set of special tokens, such as “<|user|>". 4. Scalability: We validate the effectiveness of L2T on a large-scale dataset of nearly 1M samples, whereas IM’s evaluation is limited to much smaller datasets, showing performance saturation at just 13k samples.
@@ 5 Conclusion
In this paper, we propose L2T, which enhances multimodal capabilities by learning to instruct images as a regularizer, rather than focusing only on learning to respond. This enables L2T to seamlessly expand the training data, reducing overfitting and proactively guiding the model to learn visual inputs more effectively, thereby avoiding shortcuts. Experimental results demonstrate the effectiveness of L2T across 16 multimodal tasks, highlighting its superior performance on OCR and image captioning tasks by placing greater emphasis on visual content. Furthermore, L2T significantly improves hallucination mitigation. It is also worth noting that our method is orthogonal to existing advancements in MLLMs and can be easily integrated into these methods with minimal compromises on computational costs. We believe that L2T has the potential to evolve the general VIT framework by mitigating overfitting, enhancing data efficiency, and promoting improved generalization in MLLMs.
Acknowledgement
This work is partially supported by the National Key R&D Program of China (No. 2022ZD0160703), National Natural Science Foundation of China (No. 62306178) and STCSM (No. 22DZ2229005), 111 plan (No. BP0719010). BH is supported by RGC Young Collaborative Research Grant No. C2005-24Y, RGC General Research Fund No. 12200725, and NSFC General Program No. 62376235.
@@ Appendix A Detailed Removed Task Templates
In this section, we provide a comprehensive overview of the task templates removed during the training of TinyLLaVA, LLaVA 1.5, and LLaVA-NeXT, as shown in Table˜8 and Table˜9. Notably, LLaVA-NeXT-Data includes a more diverse set of instruction datasets, leading to a wider variety of task templates. Consequently, a larger number of task templates are removed compared to those discarded from the LLaVA-mix-665k dataset.
@@ B.1 More Details on Hyperparameter
Detailed hyperparameters for training TinyLLaVA, LLaVA 1.5 and LLaVA-NeXT are shown in Table˜10. For TinyLLaVA and LLaVA 1.5, pretraining is conducted with a learning rate of 1e-3 and a batch size of 256, while finetuning uses a learning rate of 2e-5 and a batch size of 128. For LLaVA-NeXT, pretraining uses a learning rate of 1e-3 and a batch size of 128, with finetuning using a learning rate of 1e-5 and a batch size of 32. All experiments use the AdamW optimizer [Loshchilov and Hutter, 2019] and a cosine decay schedule with a warm up ratio of 0.03. During pretraining phase, only the multimodal connector is trained. In finetuning, TinyLLaVA and LLaVA 1.5 jointly train the connector and language model, while LLaVA-NeXT trains all parameters, including the vision encoder.
@@ B.2 More Details on Training Data
In the finetuning stage of instruction tuning, we use the LLaVA-mix-665k data [Liu et al., 2024a] on TinyLLaVA and LLaVA 1.5, which incorporates a diverse set of instruction-following datasets, including free conversational data (LLaVA-Instruct [Liu et al., 2023a]), visual question answering (VQAv2 [Goyal et al., 2017], GQA [Hudson and Manning, 2019], OKVQA [Marino et al., 2019], A-OKVQA [Schwenk et al., 2022]), OCR and captioning (OCRVQA [Mishra et al., 2019], TextCaps [Sidorov et al., 2020]), and visual grounding (RefCOCO [Kazemzadeh et al., 2014], VG [Krishna et al., 2017]). For LLaVA-NeXT, we use the LLaVA-NeXT-Data [Liu et al., 2024b], which extends the LLaVA-mix-665k [Liu et al., 2024a] dataset with high-quality user instruction data from LAION-GPT-V, ShareGPT-4V [Chen et al., 2024a], and LLaVA-demo [Liu et al., 2023a], as well as OCR, document, and chart data from DocVQA [Mathew et al., 2021], SynDog-EN [Kim et al., 2022], ChartQA [Masry et al., 2022], DVQA [Kafle et al., 2018], and AI2D [Kembhavi et al., 2016].
@@ Appendix C More Details on CHAIR Evaluation
As mention in Section˜3.3, we utilize the CHAIR metric to evaluate object hallucination in image captioning tasks. The metric assesses object hallucinations through two dimensions: per-instance $\text{CHAIR}_{i}$ and per-sentence $\text{CHAIR}_{s}$. The former measures the fraction of object instances that are hallucinated, while the latter determines the proportion of sentences that contain at least one hallucinated object. Definitions of these two metrics are formally provided:
@@
CHAIR ${}_{s}=\frac{|\{\text{sentences with hallucinated object}\}|}{|\{\text{ all sentences}\}|}.$
We follow [Huang et al., 2023] to perform the CHAIR evaluation on the COCO2014 dataset. Specifically, we randomly select 500 images in the validation set. The decoding process uses the greedy strategy and the Beam search with $N_{beam}=5$, with a maximum of 512 tokens for new content generated.
@@ Appendix D More Cases
More cases of our L2T compared to VIT are shown in Figure˜11. Our L2T consistently outperforms VIT across a range of scenarios, including OCR, image captioning, and hallucination mitigation.
Figure 11: More cases of L2T and VIT across OCR, image captioning, and hallucination mitigation tasks.
@@ Appendix E Computational Cost
Table˜11 presents the computational cost of VIT and L2T across different instruction-to-response length ratios. The experiments are conducted on 8 NVIDIA A100 GPUs.
Table 11: Computational cost of VIT and L2T across different instruction-to-response length ratios. We report the number of samples per batch and steps per batch. Experiments are based on LLaVA-1.5 Vicuna-7B.
@@
L2T 2.640 0.330 2.623 0.328 2.676 0.334 2.613 0.327 2.708 0.338
Appendix F Visualization of Attention Heads
Recent observations [Orgad et al., 2025, Shen et al., 2025] show that the hidden activation of the token preceding the answer (the colon “:”, in the prompt “ASSISTANT:”) encodes more information than the output logits. Inspired by this, we select this token’s hidden activation for visualization. Specifically, we visualize 14 attention heads from the last layer of a 24-layer transformer during inference, focusing on their activation across input tokens, particularly image tokens, to compare VIT and L2T. Data samples are randomly selected from VQAv2 [Goyal et al., 2017], GQA [Hudson and Manning, 2019] and OKVQA [Marino et al., 2019]. Detailed information about the selected examples is provided in Table˜12, and the visualizations of the corresponding attention heads are presented in Figure˜6 and Figure˜12.
Table 12: Examples for visualization of attention heads.
Prompt Image Visualization
[…]
(a) Case 1: Attention weights in VIT.
(b) Case 1: Attention weights in L2T.
(c) Case 2: Attention weights in VIT.
(d) Case 2: Attention weights in L2T.
(e) Case 3: Attention weights in VIT.
(f) Case 3: Attention weights in L2T.
Figure 12: Visualization of attention weights in VIT and L2T for additional cases. Darker colors indicate higher attention weights. Experiments are based on TinyLLaVA Qwen2-0.5B.
Appendix G Discussion on the Scope of This Work
Recent vision–language models (VLMs) often struggle with challenges such as hallucination and shortcut learning, which can limit their reliability and generalization. Addressing these issues requires careful consideration of the interactions between different training stages, including Supervised Fine-Tuning (SFT), rule-based Reinforcement Learning (RL), and Reinforcement Learning from Human Feedback (RLHF).
Our work focuses on improving the foundational SFT stage. We posit that the quality of early instruction tuning directly influences the effectiveness and safety of subsequent alignment. Indeed, recent studies indicate that applying RL or RLHF on models already prone to hallucination or shortcut learning can reinforce these issues, leading to confidently incorrect outputs.
L2T addresses this problem at its root by reducing shortcut learning and enhancing visual grounding during SFT. A stronger SFT foundation provides a more reliable base for RLHF, enabling safer and more effective alignment. While recent work has emphasized RL and reward modeling as post-training interventions, foundational flaws in SFT can persist or even intensify during alignment. Thus, improving early-stage instruction tuning serves as a preventive measure rather than relying solely on corrective post-training methods.
Overall, this perspective highlights the broader impact of L2T: by enhancing the base model during SFT, it facilitates downstream alignment and helps define the scope of our contribution in enabling safer and more generalizable multimodal training strategies.
Appendix H Limitations
Despite the promising capabilities underscored by L2T, several limitations must be acknowledged. First, the performance improvements achieved by L2T mainly stem from the informative instruction content that aids in better understanding visual inputs. In contrast, redundant instruction content, such as system or task templates, do not contribute to performance gains, as they lack relevant information about the visual content. Second, it is critical to ensure that the instructions do not contain harmful or biased content. The presence of such content could raise concerns about the fairness and reliability of MLLMs, making them more susceptible to generating flawed, biased, or even harmful responses.
Appendix I Broader Impact